# The Minimal Attached Eddy in Wall Turbulence: Statistical Foundations, Inverse Identification and Influence Kernels

### v4

@@ 5  Related approaches and connections
The present work occupies a complementary position within the broader landscape of attached-eddy theories and coherent-structure studies. An important parallel development is the hierarchical random additive process (HRAP) model introduced by YangMarusicMeneveau2016HRAP which provides an alternative statistical reinterpretation of Townsend’s attached eddy hypothesis. This formulation elegantly recovers logarithmic scaling laws for moments, structure functions, and generalized two-point correlations without requiring explicit specification of eddy geometry.
In this sense, the HRAP model and the marusic2019attached review reinforce a key message: the precise morphology of the prototypical eddy is not essential for reproducing log-layer statistics. The present work provides a concrete mechanistic underpinning for this observation. The inverse-problem analysis (Section 2) demonstrates that the inferred influence functions $I_{1}(\eta)$ and $I_{ij}(\eta)$ represent the essential kinematic constraints that any viable eddy must satisfy. Once these constraints are met: a near-plateau for $\eta\lesssim 1$ and appropriate decay for $\eta>1$, the detailed vortical architecture becomes secondary. The square hairpin succeeds not because it is the unique physical structure, but because it is a minimal template that satisfies these influence-function requirements while remaining consistent with Biot–Savart kinematics and wall boundary conditions. The HRAP framework abstracts this further by encoding only the statistical consequence (additive contributions from a hierarchy of scales) without committing to any particular morphology.
It is important to distinguish the present statistical/kinematic AEM from the growing body of work that seeks to identify wall-attached eddies as instantaneous, material features in DNS or experimental data. Beginning with delAlamo2006Clusters who identified self-similar vortex clusters in the logarithmic region of turbulent channels, a series of studies have extracted three-dimensional coherent structures using various detection criteria. DennisNickels2011 provided detailed experimental measurements of vortex packets in turbulent boundary layers via tomographic PIV. LozanoDuranFloresJimenez2012 characterized the three-dimensional structure of momentum-transferring Q2/Q4 events, demonstrating that attached structures become geometrically self-similar with sizes proportional to their distance from the wall. More recently, HwangSung2018 developed a clustering methodology to extract wall-attached structures of velocity fluctuations, and Cheng2019Identity; Cheng2020Uncovering employed bidimensional empirical mode decomposition to isolate attached eddies and quantify their contribution to skin-friction generation.
@@
davidson2009simple propose a deliberately simple log-layer model for streamwise fluctuations which reproduces the $k_{x}^{-1}$ behavior of the one-dimensional longitudinal spectrum in the log region and provides reasonable predictions for the variance $u^{\prime}u^{\prime}$ (up to an additive constant). A key conceptual message is that log-layer streamwise statistics are strongly constrained by (i) attachedness and (ii) scale-invariant population arguments, so one-dimensional spectra and even-order statistics can be surprisingly insensitive to the fine morphological details of the underlying structures. In that sense, the Davidson–Krogstad viewpoint supports the observation that once the mean and a small set of scaling assumptions are fixed, the streamwise spectrum in the log region has limited freedom.
@@ 6  Conclusions and Perspectives
  • Minimal Biot–Savart-consistent eddy template. A simple square hairpin-type vortex-loop template (implemented with Rankine vortex rods) together with an inviscid image system across the wall can reproduce both the inferred kernels and the corresponding mean and Reynolds-stress profiles with good accuracy (Figure 6).
  • A spectral Influence kernel. We introduce the streamwise-energy Influence kernel $I_{\phi}(\kappa_{x},\eta)$ (Section 4.1), which provides a compact representation of how a self-similar eddy footprint populates the one-dimensional spectrum as a function of scale. The spectral content of a single self-similar eddy as a function of dimensionless wavenumber and relative height is implicit in prior work, possibly dating back to  perry1982mechanism. In this work, the kernel makes the emergence of the $k_{x}^{-1}$ scaling band and its limitations explicit, and more transparent than population-level arguments alone and offers insight into constructing more sophisticated compositional eddy models. It is notable that $\int_{\kappa_{x}}I_{\phi}(\kappa_{x},\eta)d\kappa_{x}=I_{11}(\eta).$
The original goal of this work was to use the inferred influence functions to uncover an optimal eddy or packet configuration. One of the conclusions of this work is that a simple Rankine rod-based square hairpin is unusually predictive once the mean is anchored, and is able to represent attached eddy physics. The reasons are as follows:
@@
At this juncture, we remark that the AEM is a statistical and non-dynamic model, and thus prototypical eddies cannot be interpreted as literal material structures, even if they provide useful building blocks for kinematic prediction. Nevertheless, the inverse-kernel viewpoint provides a bridge between statistical requirements (what kernels are needed to match moments) and geometric hypotheses (what minimL eddy templates can realize those kernels under Biot–Savart induction). A good fit of a one-dimensional premultiplied spectrum does not uniquely identify a specific structure. Many different eddy templates can generate similar $k^{-1}$-like behavior once integrated over a scale-invariant population. Therefore, the fact that the Rankine-rod hairpin AEM matches experiments across many quantities of interest and Reynolds numbers is encouraging, but it is not, by itself, a proof that the flow is literally composed of those hairpins. Stronger discrimination can come from tests that 1D spectra smear out: two-dimensional spectra, cross-spectra, coherence, wall-normal and spanwise components, vorticity-related spectra, and conditional structure signatures.
@@
The inverse formulation suggests several immediate extensions: (i) simultaneous inference of multiple kernels using multiple moments and multiple Reynolds numbers with shared regularization and physically motivated constraints, (ii) inference of 2D spectral kernels (in $(k_{x},k_{z})$) rather than only 1D kernels, and (iii) incorporation of non-Poisson population effects (packets, clustering, exclusion) through pair-correlation corrections to Eq. (11) and their spectral counterparts.
Acknowledgement This research was supported by NASA grant # 80NSSC18M0149 (Technical monitor: Dr. Gary Coleman).
Appendix A: Numerical convergence evidence for the square hairpin + image.
To verify the decay/convergence directly, we performed a Biot–Savart calculation for a simple “square hairpin” U-shape (three straight segments with unit circulation) together with an opposite-sign mirror image across $y=0$, consistent with the construction described around Fig. 5. A small vortex-core regularization was used to smooth the near-filament singularity; this does not affect the far-field decay that controls convergence of the planar integrals.
@@
Direct growth of $I_{1}$ with domain size. For the same hairpin at $\eta=0.5$, integrating over $R,R]^{2}$ yields: The observed approach is consistent with the $\mathcal{O}(1/R)$ tail implied by (22). In particular, extrapolating with the model $I_{1}(\eta;R)=I_{1,\infty}(\eta)+C/R$ gives
@@
Table 1: Convergence of the planar influence integral $I_{1}(\eta;R)$ for the square hairpin + image at $\eta=0.5$.

### v5

@@ 6  Related approaches and connections
The present work occupies a complementary position within the broader landscape of attached-eddy theories and coherent-structure studies. An important parallel development is the hierarchical random additive process (HRAP) model introduced by Yang et al. (2016) which provides an alternative statistical reinterpretation of Townsend’s attached eddy hypothesis. This formulation elegantly recovers logarithmic scaling laws for moments, structure functions, and generalized two-point correlations without requiring explicit specification of eddy geometry.
In this sense, the HRAP model and the Marusic & Monty (2019) review reinforce a key message: the precise morphology of the prototypical eddy is not essential for reproducing log-layer statistics. The present work provides a concrete mechanistic underpinning for this observation. The inverse-problem analysis (Section 2) demonstrates that the inferred influence functions $I_{1}(\eta)$ and $I_{ij}(\eta)$ represent the essential kinematic constraints that any viable eddy must satisfy. Once these constraints are met: a near-plateau for $\eta\lesssim 1$ and appropriate decay for $\eta>1$, the detailed vortical architecture becomes secondary. The rectangular hairpin succeeds not because it is the unique physical structure, but because it is a minimal template that satisfies these influence-function requirements while remaining consistent with Biot–Savart kinematics and wall boundary conditions. The HRAP framework abstracts this further by encoding only the statistical consequence (additive contributions from a hierarchy of scales) without committing to any particular morphology.
It is important to distinguish the present statistical/kinematic AEM from the growing body of work that seeks to identify wall-attached eddies as instantaneous, material features in DNS or experimental data. Beginning with del Álamo et al. (2006) who identified self-similar vortex clusters in the logarithmic region of turbulent channels, a series of studies have extracted three-dimensional coherent structures using various detection criteria. Dennis & Nickels (2011) provided detailed experimental measurements of vortex packets in turbulent boundary layers via tomographic PIV. Lozano-Durán et al. (2012) characterized the three-dimensional structure of momentum-transferring Q2/Q4 events, demonstrating that attached structures become geometrically self-similar with sizes proportional to their distance from the wall. More recently, Hwang & Sung (2018) developed a clustering methodology to extract wall-attached structures of velocity fluctuations, and Cheng et al. (2019, 2020) employed bidimensional empirical mode decomposition to isolate attached eddies and quantify their contribution to skin-friction generation. Additionally, several studies have demonstrated that statistical properties can be recovered from stochastic representations of eddies with wall-distance scaling (Cuevas Bautista et al., 2019; Ehsani et al., 2024; Nath & Hickey, 2026), and the clustering and larger-scale organization of attached eddies has been investigated by Hutchins et al. (2012) and Deshpande et al. (2023), among others.
@@
We also note a connection to the resolvent framework (McKeon, 2019): the self-similar hierarchy of response modes identified by resolvent analysis shares the same wall-attached scaling as the AEM, and the resolvent modes can be viewed as a dynamically informed basis for the eddies that the AEM treats kinematically. Integrating the spectral influence kernel $I_{\phi}$ with resolvent-based predictions of the two-dimensional spectral tensor represents a promising path toward a dynamically grounded attached-eddy model.
Davidson & Krogstad (2009) propose a deliberately simple log-layer model for streamwise fluctuations which reproduces the $k_{x}^{-1}$ behavior of the one-dimensional longitudinal spectrum in the log region and provides reasonable predictions for the variance $u^{\prime}u^{\prime}$ (up to an additive constant). The existence and criteria for a $k_{x}^{-1}$ spectral dependence have been investigated extensively, more recently by Nickels et al. (2005). Notably, the $k_{x}^{-1}$ trend is most apparent very close to the surface, whereas the present comparisons are shown at an intermediate wall-normal distance. A key conceptual message is that log-layer streamwise statistics are strongly constrained by (i) attachedness and (ii) scale-invariant population arguments, so one-dimensional spectra and even-order statistics can be surprisingly insensitive to the fine morphological details of the underlying structures. In that sense, the Davidson–Krogstad viewpoint supports the observation that once the mean and a small set of scaling assumptions are fixed, the streamwise spectrum in the log region has limited freedom.
@@ 7  Conclusions and Perspectives
  • Minimal Biot–Savart-consistent eddy template. A simple rectangular hairpin-type vortex-loop template (implemented with Rankine vortex rods) together with an inviscid image system across the wall can reproduce both the inferred kernels and the corresponding mean and Reynolds-stress profiles with good accuracy (Figure 6).
  • A spectral Influence kernel. We introduce the streamwise-energy Influence kernel $I_{\phi}(\kappa_{x},\eta)$ (Section 4.1), which provides a compact representation of how a self-similar eddy footprint populates the one-dimensional spectrum as a function of scale. The spectral content of a single self-similar eddy as a function of dimensionless wavenumber and relative height is implicit in prior work, possibly dating back to  Perry & Chong (1982). In this work, the kernel makes the emergence of the $k_{x}^{-1}$ scaling band and its limitations explicit, and more transparent than population-level arguments alone and offers insight into constructing more sophisticated compositional eddy models. It is notable that $\int_{0}^{\infty}I_{\phi}(\kappa_{x},\eta)d\kappa_{x}=I_{11}(\eta).$
  • Exact closed-form analytical results for the hairpin family. For the three-segment straight-filament hairpin family with wall image, we derive exact expressions for the mean kernel $I_{1}^{\mathrm{hp}}(\eta)=\Gamma[b_{0}+(b_{1}-b_{0})\eta]H(1-\eta)$ (Appendix C, Eq. (22)), demonstrating that only the rectangular hairpin ($b_{1}=b_{0}$) produces an exact plateau and hence a purely logarithmic mean profile. Closed-form Fourier-space expressions for the streamwise velocity (Eq. (40)) enable an exact head/leg/cross decomposition of the spectral kernel, quantifying the mean-variance duality: the head contributes all of $I_{1}$ but only $\sim 17\%$ of $I_{\phi}$ at $\kappa_{x}=0$ for $\eta=0.5$, while the legs carry $\sim 53\%$ of the spectral energy.
  • Scale-dependent eddy density. Allowing $\beta$ to vary with eddy size $h$ provides an additional degree of freedom that yields near-perfect predictions of both mean velocity and streamwise variance across $Re_{\tau}=6000$2000020000 (Figure 19), with a simple piecewise-linear analytical fit (Eq. (28)).
The original goal of this work was to use the inferred influence functions to uncover an optimal eddy or packet configuration. One of the conclusions of this work is that a simple Rankine rod-based rectangular hairpin is unusually predictive once the mean is anchored, and is able to represent attached eddy physics. The reasons are as follows:
@@
At this juncture, we remark that the AEM is a statistical and non-dynamic model, and thus prototypical eddies cannot be interpreted as literal material structures, even if they provide useful building blocks for kinematic prediction. Nevertheless, the inverse-kernel viewpoint provides a bridge between statistical requirements (what kernels are needed to match moments) and geometric hypotheses (what minimal eddy templates can realize those kernels under Biot–Savart induction). A good fit of a one-dimensional premultiplied spectrum does not uniquely identify a specific structure. Many different eddy templates can generate similar $k^{-1}$-like behavior once integrated over a scale-invariant population. Therefore, the fact that the Rankine-rod hairpin AEM matches experiments across many quantities of interest and Reynolds numbers is encouraging, but it is not, by itself, a proof that the flow is literally composed of those hairpins. Stronger discrimination can come from tests that 1D spectra smear out: two-dimensional spectra, cross-spectra, coherence, wall-normal and spanwise components, vorticity-related spectra, and conditional structure signatures.
@@
The inverse formulation suggests several immediate extensions: (i) simultaneous inference of multiple kernels using multiple moments and multiple Reynolds numbers with shared regularization and physically motivated constraints, (ii) inference of 2D spectral kernels (in $(k_{x},k_{z})$) rather than only 1D kernels, and (iii) incorporation of non-Poisson population effects (packets, clustering, exclusion) through pair-correlation corrections to Eq. (11) and their spectral counterparts. A key limitation of the present validation is that it is restricted to the streamwise velocity component: the mean $U^{+}$, the streamwise variance $\overline{u^{\prime 2}}$, and the 1D streamwise energy spectrum. Extending the comparisons to all components of the Reynolds stress tensor (wall-normal and spanwise variances and the Reynolds shear stress), two-point correlations, and two-dimensional $(k_{x},k_{z})$ spectra would provide a much more stringent test of the minimal eddy hypothesis and could distinguish between eddy templates that appear equivalent in 1D metrics.
Acknowledgement This work was supported by the Aeronautics Research Mission Directorate at NASA through the Transformative Aeronautical Concepts Program (TACP) and the D.2 Transformational Tools and Technologies Project (TTT) under contract # 80NSSC23M0215. Early part of the research was supported by NASA grant # 80NSSC18M0149. The analytical expressions in Equations 36-39 were derived with the help of GPT5.2-Pro from OpenAI, Inc.
Appendix A: Numerical convergence
$\int_{\mathbb{R}^{2}}|u_{1}(\xi,\zeta,\eta)|\,d\xi\,d\zeta<\infty.$ (29)
Reynolds-stress kernels are less restrictive.
Similarly, the second-moment influence functions are
For line-vortex Biot–Savart fields, a generic open filament can generate a $\rho^{-2}$ contribution to certain velocity components in the far field, which is precisely the borderline case for $I_{1}$ and would lead to a logarithmic growth of $I_{1}(\eta;R)$ with $R$. The specific square-hairpin eddy used herein is paired with an opposite-sign mirror image across $y=0$ (to enforce inviscid wall boundary conditions), and the wall-parallel segment cancels with its image. This construction removes leading-order contributions responsible for a $\rho^{-2}$ tail in the streamwise component and yields a faster far-field decay. In other words, the image construction is not a technical convenience: it is central to ensuring that Eq. (4) is well posed for a line-vortex eddy model.
Numerical evidence.
To verify the decay/convergence directly, we performed a Biot–Savart calculation for a simple rectangular hairpin (three straight segments with unit circulation) together with an opposite-sign mirror image across $y=0$, consistent with the construction described around Fig. 5. A small vortex-core regularization was used to smooth the near-filament singularity; this does not affect the far-field decay that controls convergence of the planar integrals.
@@
Direct growth of $I_{1}$ with domain size. For the same hairpin at $\eta=0.5$, integrating over $R,R]^{2}$ yields the values shown in Table 1. The observed approach is consistent with the $\mathcal{O}(1/R)$ tail implied by (32). In particular, extrapolating with the model $I_{1}(\eta;R)=I_{1,\infty}(\eta)+C/R$ gives
@@
Table 1: Convergence of the planar influence integral $I_{1}(\eta;R)$ for the rectangular hairpin + image at $\eta=0.5$.